T.R | Title | User | Personal Name | Date | Lines |
---|
61.6 | Table Stpping Problem after polish. /Step #6 value incorrect | ASDG::POIRIER | | Fri Nov 10 1995 20:04 | 9 |
|
CMP.A2 Table Stopping Condition:
Eric Bodensieck found the cause of the problem. It seems that Step #6
had a value of 1 sec entered for the Table. Please remember that all
values in the step #6 screen should be zero.
John.
|
61.7 | Reworked Speedfam Carrier Rings | ASDG::POIRIER | | Thu Nov 30 1995 19:40 | 11 |
|
Reworked Carrier rings from Speedfam:
The rings have arrived and are in the carrier rebuild room.
I tried working with 3 of them and all 3 had delta exposures of
.002 or more after being tightened. D shift, see if you have
better luck. Maybe the shims need deburring?
Also, I left 16 M4 screws with lock & flat washers on the table.
Use these fastners to on 2 of the carriers per Tim Dyer. We want to
see if the loosening problem decreases with this set-up.
|
61.8 | CMP Carrier Status | ASDG::POIRIER | | Sat Dec 16 1995 19:05 | 26 |
|
* Carrier Status:
--------------
Strasbaugh Carriers: 4 completed Carriers in the Vidmar Cabinet.
(2 built by Leon, 1 by John & 1 by Brad)
2 Filmed & Shimmed Plates in the Vidmar.
(1 by John & 1 by Brad)
2 experimental Carriers on the Storage Rack
with R-200 film. Please hold for Kyle.
(2 built by John)
Speedfam Carriers:
-----------------
10 completed Carriers are on the storage rack.
We are still waiting for the backordered rings.
Tim Dyer needs 2 more used Carrier Plates for
his Speedfam R-200 film experiment. Please place
the next 2 used plates on the bottom shelf
of the storage rack for Tim.
|
61.1 | DRILL BITS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Dec 20 1995 21:36 | 10 |
| ASDG::POIRIER 7 lines 14-SEP-1994 23:33
-< New drill bits in tool box. >-
--------------------------------------------------------------------------------
* We now have #7 and #21 drill bits in the tool box. Use the #21 drill
bit with the 10-32 tap and the #7 drill bit when tapping 1/4-20 holes.
John.
|
61.2 | Strasbaugh Cond. Disk & Carrier weight values | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Dec 20 1995 21:36 | 19 |
| ASDG::POIRIER 16 lines 28-JAN-1995 19:06
-< Strasbaugh Cond. Disk & Carrier weight values. >-
--------------------------------------------------------------------------------
* Please note and save the following weights. Use these values when
performing Spindle Down Force Calibrations and Conditioner Down
Force Calibrations on the Strasbaugh Polishers.
Load Cell Carrier: 7.73 lbs. lbs.
Process Carrier: 22.35 lbs.
4" Conditioning Disk: 1.18 lbs.
6" Conditioning Disk: 1.83 lbs.
John.
|
61.3 | Instruments on a set calibration schedule | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Dec 20 1995 21:37 | 46 |
| SUBPAC::LANDRY 43 lines 25-MAR-1995 10:50
-< Instruments on a set calibration schedule >-
--------------------------------------------------------------------------------
Description Cal # S/N or M/N Manufacturer
--------------------------------------------------------------------------------
80TK Thermocouple Module FS164 M = 80TK Fluke
80TK Thermocouple Module FS165 M = 80TK Fluke
80TK Thermocouple Module FS193 M = 80TK Fluke
80TK Thermocouple Module FS197 M = 80TK Fluke
Fluke 27 Multimeter FS163 S = 5905038 Fluke
Fluke 27 Multimeter FS196 S = 5905066 Fluke
Fluke 87 Multimeter FS162 S = 59170240 Fluke
Fluke 87 Multimeter FS161 S = 59170241 Fluke
Current Transformer FS177 M = 80I-600 Fluke
Torque Wrench FS178 S = 12930010714 Snap-on
Torque Wrench FS205 S = 09940001869 Snap-on
Torque Wrench FS203 S = 09940001790 Snap-on
Torque Wrench FS202 S = 09940001947 Snap-on
Torque Wrench FS200 S = 9-94 1108 Snap-on
Torque Wrench FS201 S = 9-94 1130 Snap-on
Torque Wrench FS220 S = 0301 Snap-on
Torque Meter FS219 S = 114048 Snap-on
Dial Indicator (carriers) FS198 S = 009922 Mitutoyo
Dial Indicator (carriers) FS212 S = 009924 Mitutoyo
Tachometer FS156 S = L227190 Extech Instr.
Tachometer FS194 S = L227098 Extech Instr.
Caliper FS208 S = 7182054 Mitutoyo
Caliper FS207 S = 7021463 Mitutoyo
Caliper FS155 M = 120 Starrett
Depth Indicator FS154 M = 445 Starrett
Depth Indicator FS192 M = 445 Starrett
Dial Indicator FS150 M = 711 Starrett
Dial Indicator FS191 M = 711 Starrett
Dial Indicator FS151 M = 25-111 Starrett
Dial Indicator FS152 M = 25-441 Starrett
Dial Indicator FS153 M = 25-131 Starrett
Micrometer FS149 M = 436-1 Starrett
Micrometer FS190 M = 436-1 Starrett
Flatness Gauge FS213 M = J1K Crane Pack. Co.
Infrared Monitor FS160 M = OS52 Omega
Infrared Monitor FS195 M = OS52 Omega
Pressure Indicator/Load Cell FS167 S = 252004 Sensotec
Pressure Indicator/Load Cell FS214 S = 325137 Sensotec
|
61.4 | Activities & Events on the Strasbaughs since 4/5/95 | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Dec 20 1995 21:38 | 77 |
| ASDG::POIRIER 78 lines 19-APR-1995 14:16
-< Activities & Events on the Strasbaughs since 4/5/95 >-
CMP.A1
------ Date
----
1) Table stalled, Robot End Effector vacuum
failure problem resolved. 4/5
2) Carriers leaving wafers on Table at end
of Auto Cycle. Weekly PM completed.
Terminal strip moved to new location. 4/7
3) POU filters installed. Swapped DC Servo
Module #2 with #4. 4/10
4) New DC Servo Module installed for Axis #2. 4/12
5) Started Verification run. Replaced Tygon tubing
on Slurry pump #1. Calibrated slurry flow. 4/13
6) Weekly PM completed. 4/14
7) Verification run completed. Table, Spindle & Hub
alignments checked. (OK) Uneven backpressure readings
noticed on 2nd shift. Carriers removed and o-rings replaced.
Water drained from filters in Bridge. Backpressure stablized.
New conditioning disk installed. 4/18
8) Backpressure calibration checked. (OK)
Table, Right Spindle & Left Spindle RPM checked. (OK)
New carriers built (with curved plates). Waiting for
uniformity results. 4/19
CMP.A2 Date
------ ----
1) Moved Romote to WIP isle due to monitor
failure. Swapped cables. Carriers sticking
to Pad. Carrier removed and rebuilt. 4/5
2) System failed to finish an auto cycle after
a FINISH CYCLE command was entered.
Weekly PM started. Replaced Carriers
due to poor uniformity. 4/7
3) Installed POU slurry filters. 4/8
4) Weekly PM completed. 4/10
5) Broke wakers on Pad. Carriers disassembled. (OK)
Ashland replaced valve in slurry distribution box
causing pressure to increase and bypass peristaltic
pumps. Slurry POU filters clogged and were replaced.
Table stalled for the first time. 4/13
6) DC Servo Module for axis #4 replaced. 4/17
7) Slurry POU filters clogged and were removed.
Bypass installed. New conditioning disk installed. 4/18
8) Weekly P.M. completed. Table, Spindle & Hub
aligment checked. (One Spindle was slightly out).0035"
Pad replaced with an (IC-1000 SUBA backed) Speedfam Pad.
Waiting for uniformity results using the new Pad. 4/19
Carrier consumables:
-------------------
Check valves - 9 new check valves from the bin on the
Carrier rebuild bench were checked.
8 were O.K., 1 failed. 4/19
|
61.5 | New Load Cell 0 - 50 lbs | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Dec 20 1995 21:39 | 14 |
| ASDG::POIRIER 11 lines 14-JUN-1995 19:23
-< New Load Cell 0 - 50 lbs. >-
--------------------------------------------------------------------------------
* Load Cell for measuring Conditioner down force:
The new SENSOTEC 0 - 50 lb. Load Cell arrived today.
It will be used for calibrating the Conditioner down
force on both Strasbaugh Polishers. It has been labled
(CONDITIONER DOWN FORCE ONLY) and is stored in the grey
roll around cabinet in the service chase. Tim is adding
it to the calibration list for us.
John.
|
61.9 | Cmp.b2 index table | ASDG::POIRIER | | Thu Dec 21 1995 19:32 | 7 |
|
* CMP.B2 Index Table:
The (old) damaged index table from cmp.b2 was returned to speedfam
for warranty credit.
J.P.
|
61.10 | OnTrak OCP's repaired and back on-line. / Jumpers removed. | ASDG::POIRIER | | Fri Dec 22 1995 17:00 | 12 |
|
2% Amonium Hydroxide OCP's:
--------------------------
The OCP's that monitor Amonium Hydroxide for both scrubbers have been
repaired by Ashland. They finally traced the problem to a bad EPROM
in the Systems Chemistry Unit down in the sub fab.
The jumpers have been removed from both OnTrak electrical cabinets.
The Scrubbers will now respond to a NH40H pressure detect from the
OCP's. Both OCP's are now in the normal "DISP" state.
|
61.12 | Start-up for 12/29 | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 29 1995 15:12 | 7 |
|
For start-up.... there are (4) Strasbaugh carriers
ready to go. There are also (17) SpeedFam carriers ready
as well.....
BJ
|
61.13 | CMP Start-up Activities / High Putrity D.I. Down | ASDG::POIRIER | | Sat Dec 30 1995 15:51 | 18 |
|
* CMP Start-up Status:
-------------------
Mitch & I could not proceed with the start-up activities due to
the High Purity D.I. Water being out of spec. The latest
communication from the CFM Group is that the D.I. Water is still
out of spec as of 3:00 p.m. Saturday. It won't be up until 7:00 Sunday
morning at the earliest.
The CFM Group contact performing the D.I. Water analysis is Diana
Roberts. She can be paged at 508 489-6859. Mitch will be in Sunday
morning for the latest update on the D.I. Water. If the water comes
up, he'll call me at home and I'll come in. If any of you CMP EE
techs are reading this message from home and want to come in on
Sunday, call Mitch in the fab or page him at 652.
J.P.
|
61.14 | pressure plates...... | SUBPAC::LANDRY | | Wed Jan 03 1996 19:43 | 19 |
| The last batch of pressure plates received in from Speedfam have showed
that the spec'd 2 um cirvature have indeed been way out. Some have
ranged up towards 20 um, resulting in poor uniformity. They have been
removed from the polishers and are being sent back to Speedfam for
refilming.
15 plates were sent back this afternoon.
10 were sent back the day before.
There are some more that were removed from the system late today that
will need to go back as well. NO SBA has been cut for these as of yet.
PLEASE read Frank Krupas message about saving the data that accompanies
the shipment of new pressure plates. They reflect the micron
cirvatures that we receive.
Doc
|
61.15 | Speedfam Carrier Plates | ASDG::POIRIER | | Thu Jan 04 1996 19:00 | 6 |
|
* Speedfam Carrier Plates:
-----------------------
5 Speedfam Carrier Plates arrived at 10:30 A.M. this morning.
They were used to build 5 Carriers which are now running on CMP.B2.
|
61.16 | Speedfam Carrier Plates | ASDG::POIRIER | | Sat Jan 06 1996 19:36 | 9 |
|
* Speedfam Plates:
---------------
10 Speedfam Plates arrived today. 5 were built and 5 still need
building. I removed 10 Speedfam Carrier Rings from the Fab6
Stockroom today. There are 19 left in stock.
|
61.17 | plates | SUBPAC::LANDRY | | Tue Jan 09 1996 19:28 | 13 |
| * I have generated an SBA to send back 5 Speedfam pressure plates.
It's in the loop, waiting approval. I'll bring the plates to the
shipping receiving dock tomorrow.
* Please be sure to keep the good plates seperated from the bad. The
good plates should be stored in the corner, next to the rebuilding
table. The bad ones should be put back into the transport holders,
then placed into the case and left next to the Vidmar cabinet.
* Opts has built all available plates/carriers for the Speedfam (8).
Doc
|
61.18 | DON'T GREASE C. ARM BEARING. | FABSIX::B_WESTPHAL | | Fri Jan 12 1996 19:45 | 3 |
| Tim requested that the Speedfam Conditioning arm gimble bearing no
longer be lubed with grease, it is to be left dry. This is what was
recommended by the speedfam reps.
|
61.19 | SBA's X 3 = much lugging around | SUBPAC::LANDRY | | Thu Jan 18 1996 12:35 | 16 |
| Couple of quick updates..........
* Shipped out three Strasbaugh plates to the machine shop for repair.
All have broken bolts or taps in the jack screw threads.
* Shipped out 5 more pressure plates to Speedfam. These plates were
found squirreled away in the office area. We should be careful not
to let this happen as we need to have 20 plates in-house at all
times. We only have a flow loop of 35 plates.
* Shipped back all the replaced parts from B1's multihead that were
removed this week (solenoids, gauges, rotary unions, vacuum tank,
scrubber bracket, etc.)
Doc
|
61.20 | assorted shipment news, lost cal equip | SUBPAC::LANDRY | | Mon Jan 22 1996 18:26 | 21 |
| Today's generic activity..............
* Received in 10 Speedfam pressure plates. They are in cases in the
carrier rebuild room.
* Shipped out 7 used pressure plates to Speedfam for refilming.
* Shipped out the broken Index Table load cup (from B2), a bad
Infranor power supply and a bad Infranor controller to Speedfam.
* Looked extensively for (but could not locate) a Sensotec force
measurement gauge. It is the original down force measurement
gauge for the Speedfam. If anyone knows of it's location, please
let a first shift EE tech know, so that we may submit it for cal.
It has a Digital calibration sticker with the # FS218. As an FYI,
it is NOT the gauge that is in the black case by the Speedfam
cabinet. That one will be added to the cal list once the proper
credentials arrive.
A shift
|
61.22 | SLurry system failure / send carrier plates to speedfam | ASDG::POIRIER | | Sat Feb 17 1996 19:40 | 25 |
|
* Speedfam Pressure Plates need to be sent back to Speedfam for film
replacement. There is a box of 5 ready to go in the carrier rebuild
room. You may want to box-up another 5 due to tonight's crash.
A shift, please try to ship them out on Monday.
C Shift.
P.S. A Shift, we sent out the 2 boxes that you requested on Friday.
* Slurry system failure:
---------------------
The Slurry system alarm went off at 7:35 this evening. B Shift,
call Ashland and make sure noone runs any polishers until the
slurry pressure is back to normal.
ASHLAND 24 HR ON-CALL PAGER # 508-489-4393
|
61.23 | Ammonium hydroxide down, Slurry System down. | ASDG::POIRIER | | Sat Feb 17 1996 20:06 | 8 |
|
* Scrubbers down, Polishers down:
------------------------------
The nh40h for the scrubbers & the slurry for the polishers are
inactive. This event happened at shift change and "B" shift is
in the process of contacting Ashland Chemical for assistance.
C Shift.
|
61.24 | Slurry & Ammonia Shutdown | FABSIX::R_GEE | | Sun Feb 18 1996 05:24 | 10 |
| * At the end of "C" shift the slurry and ammonia shutdown.
We called Ashand Chemical and they informed use that
somebody in the fab had pushed one of the EMO's shutting
down all the chemical delivery systems in fab6. Found the EMO
in bay 6.2n 2 had been pushed for no apparent reason.
Unlocked the EMO and Ashland's reset the chemical delivery
systems.
B Shift
|
61.25 | 28" Rodel IC-1000 Pads performing Poorly | ASDG::POIRIER | | Sat Feb 24 1996 19:32 | 12 |
|
* 28" Rodel Pads on the Strasbaugh's:
-----------------------------------
FYI:
Manufacturing has been struggling with the Pads all day. It appears
the the Pads will not perform to spec if they have been on the
Polishers for 24 hrs. This happens even if the wafer count is under
100. The tools sat all night, A1 had 98 wafers on it and A2 had 135.
Both Pads would not pass MQC's. It took 50 monitor wafers to break-in
the Pad on A2.
|
61.26 | NO AMMONIA FLOW | FABSIX::R_GEE | | Sun Mar 03 1996 07:44 | 3 |
| AN AMMONIA TO THE ROOM, ASHLAND'S BEEN NOTIFIED
B SHIFT
|
61.27 | micrometer missing?? | FABSIX::B_FINN | | Mon Mar 04 1996 19:24 | 34 |
| Greetings:
The following items have been submitted to the Cal Lab
for calibration, on March 4, 1996.
LANDRY TIM FS0150 DIAL INDICATOR STARR 711 5/13/96
LANDRY TIM FS0155 Dial Caliper 12" STARR 120-12 5/03/96
LANDRY TIM FS0167 PORTABLE INDICATOR SENSO 060-3155-01 5/08/96
LANDRY TIM FS0191 DIAL INDICATOR STARR 711 FS191 4/19/96
LANDRY TIM FS0198 DIGIMATIC INDICATOR MITUT 543-611B 009922
LANDRY TIM FS0206 CALIPER 24 IN. MITUT 500-506 3/12/96
Owner Notes: spare per tim landry 7/13/95
LANDRY TIM FS0263 DIGIMATIC INDICATOR MITUT 543-611B 01353 5/03/96
LANDRY TIM FS0265 Electronic Micrometer STARR 732 5/07/96
The following micrometer could not be located. If you have it or know
of its where abouts please submit it to the cal lab and place a
message here.
***** LANDRY TIM FS0190 MICROMETER, 1 IN. STARR 436-1 4/19/96 ******
A shift
|
61.28 | Speedfam Index Table Return | FABSIX::B_FINN | | Mon Mar 04 1996 19:36 | 5 |
|
Moved INDEX TABLE and create to fab six dock, along with SBA for
Shipment back to Speedfam.
A shift
|
61.29 | Solvent Sink Faucet leak / CMP.A! TCU Temp / Please CLean Carriers | ASDG::POIRIER | | Sat Mar 09 1996 19:41 | 25 |
|
* Solvent Sink:
------------
The faucet is starting to leak more severly. The top of the sink deck
now has a permenamnt puddle. It will need attention soon.
* CMP.A1 TCU:
----------
The Temp on CMP.A1's TCU has been returned to its normal setting.
It was temporaraly set to 35 deg C for experimentation purposes.
We returned the setting to 37 deg C Today.
* FYI: Please remember to bring all slurry laden carriers to the
stainless topped table in the chase for cleaning and disassembly
BEFORE transporting them to the Carrier Rebuild Room.
If there isn't time to clean & disassemble the carriers, place
them in a tray of D.I. Water. Five slurry encrusted carriers
were found in the Carrier Rebuild Room. The clean-up activities
lasted for 1 1/2 hours due to the dried slurry condition.
Thanks,
C Shift.
---
|
61.30 | plates...... | SUBPAC::LANDRY | | Tue Mar 12 1996 12:45 | 6 |
| Today.......
WE shipped out 10 Speedfam plates for re-filming.
A Shift
|
61.31 | Strasbaugh BackPressure Recipe Changes | ASDG::POIRIER | ASDG::POIRIER | Thu Mar 21 1996 20:17 | 17 |
|
* STRASBAUGH BACKPRESSURE RECIPE'S:
---------------------------------
Please be aware that all recipies on CMP.A1 & CMP.A2 have been
modified. There is no backpressure present in any recipe.
The 2 Carriers currently on CMP.A1 and the 2 on CMP.A2 have had
their check valves & White Silicon o-rings removed. This means
they are now Single Zone Carriers. Do not apply ANY backpressure
while running these Carriers.
* A work order for a telephone in the Carrier Rebuild Room has
been approved by Kim Conway.
* An SBA was generated for 10 Carrier Plates to be returned to
Speedfam for film conturing.
|
61.32 | slurry supply "qlitch" | SUBPAC::LANDRY | | Wed Mar 27 1996 18:35 | 11 |
|
The slurry supply system kicked off momentarily today. The
contractors working in the room activated a proximatey switch
on the slurry VMB's. Ashland people, fortunately, were in the
slurry room as the mishap happened. They reset the pump as
quick as humanly possible. It's possible the major vibration
that occured with B2 around the same time of day was caused by
this loss of slurry. See B2's note for details.
A Shift
|
61.33 | 10 Speedfam Plates returned for Film Conturing | ASDG::POIRIER | | Thu Mar 28 1996 19:45 | 5 |
|
Speedfam Pressure Plates:
------------------------
10 Pressure Plates were returned to Strasbaugh for Film Conturing Today.
|
61.34 | High Purity D.I. Water Pipe Break | ASDG::POIRIER | | Sat Apr 13 1996 20:33 | 11 |
|
* High Purity D.I. Water Pipe Break:
A contractor working in the new CMP bay accidentally stepped
on one of the High Purity D.I. water Taps at 11:30 a.m.
Water gushed out of the main for approx 20 min before it
could be shut off. The tap connection was repaired and water press
was restored around 1:00 p.m. The resistivity of the D.I. water
was verified at 2:45 p.m. The scrubbers are back up and MQC'd, CMP.A1
is running break-ins & CMP.A2, B1 & B2 are running MQC's.
|
61.35 | power sag | SUBPAC::LANDRY | | Tue May 07 1996 20:33 | 9 |
|
The electrical power in the room sagged today. The room lights
dimmed for a second. CMP.A1 & CMP.A2 both glitched. A1's screen
was "Windows Hyrogliphics" and A2's screen dumped out to DOS. A
quick reboot and homing took care of the beasts. No other tools
were affected by the sag.
Doc
|
61.36 | Slurry filter clogged | FABSIX::R_GEE | | Tue May 21 1996 05:18 | 5 |
| The room was down from 2:45am until 4:00am due to
no slurry flow. Called in Ashland Chemical to replace
a clogged slurry filter in the subfab.
B Shift
|
61.37 | RHS bridge drive belts | SUBPAC::LANDRY | | Tue May 21 1996 15:57 | 6 |
| Two drive belts, used to move the Strasbaugh bridge, were received
in today. They are currently located in the file cabinet outside
of James Tooker's office.
Doc
|
61.38 | Speedfam software memo | SUBPAC::LANDRY | | Tue May 21 1996 18:10 | 8 |
|
ALL THREE SPEEDFAM POLISHERS ARE NOW ON THE SAME REV 493C SOFTWARE. PLEASE
REPORT ANY BUGS IN THE NOTES FILE UNDER "BUGS/WISH". DO NOT CHANGE THE
REV SOFTWARE ON THE TOOLS WITH OUT CONTACTING MYSELF.
Thanks,
Tim Dyer
|
61.39 | Slurry pressure | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Jun 01 1996 05:21 | 7 |
|
Slurry pressure in the room was WAY TOO HIGH. A2 was
registering 45 psi, slurry was dripping out on the table
during WET IDLE and has high particles. A1 blew a slurry
line. Used by-pass by B3 and opened it a 1/4".
BJ
|
61.40 | Room down due to no slurry | FABSIX::R_GEE | | Mon Jun 03 1996 08:18 | 5 |
| The room was down from 5:00am until 7:00am due to
no slurry flow. Called in Ashland Chemical to replace
a clogged slurry filters in the subfab.
B Shift
|
61.41 | load cell shimmed | FABSIX::J_SWENSON | | Thu Jul 11 1996 20:30 | 6 |
| Speedfam downforce calibration fixture has been shimmed.
The load cell has been shimmed into it's counter bored hole so
it will now fit snug. Whom ever uses it next, please observe any
improvements/problems with the performance of the fixture.
c-shift
|
61.42 | plates / adhesives | SUBPAC::LANDRY | | Wed Jul 17 1996 20:40 | 10 |
| Today.....
* brought up some Speedfam pressure plates from the dock
* shipped out one Strasbaugh plate to the machine shop for repair
* ordered some Strasbaugh "double sided" adhesive for the brushes
* submitted a SID sheet to stock above mentioned adhesive as a
separate entity. This will allow us to get away from the 1_to_1
ratio of brushes:adhesives. I will post the P/N when they arrive.
A shift
|
61.43 | C shift please....... | SUBPAC::LANDRY | | Tue Jul 23 1996 20:41 | 11 |
| A couple of "need done" items.......
* On Bob Gee's desk, lies a a box containing a Strasbaugh carrier.
This needs to be brought to the carrier rebuild room. It's back
from repair.
* There are some Speedfam pressure plates on the Fab 6 dock, "C shift,
Please" pick them up as the off shift lads can't access that area.
A shift
|
61.44 | adhesive backings for the RHS brushes | SUBPAC::LANDRY | | Tue Jul 30 1996 10:08 | 22 |
|
Greetings all,
As most of you know, we have a discrepancy between the the number
of Strasbaugh scrubber brushes and the adhesive supplied with the
brush. Due to the 1-to-1 stocking of these two, if an adhesive is
put down wrong, it's necessary to take another adhesive backing and
re-apply it. This, naturally, throws off the inventory status in
stockroom 6. The brush vendor (Cupp Industries) supplies the backing
with the brush.
So, what is being done to offset this problem, you may ask? A one
time purchase of adhesive will take place. It will be given to us
to store in the fab (NOT the stockroom). It will be there for emergency
back up should you need it. There is a minor draw-back to this buy.
We will be receiving a roll of tape (NOT the pre-cut pieces we receive
with the brushes). The width of the roll will match the circumference
of the brush. It will be necessary to cut the roll, much like we cut
the ID of a polish pad, to match the mounting fixture. A small incon-
vienience.
Doc
|
61.45 | Slurry system shutdown | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Aug 08 1996 07:53 | 7 |
|
The slurry system kicked off at 5am because of
clogged filters downstairs. It finally came back
on at 6:15am when someone from Ashland showed up.
BJ
|
61.46 | Chemicals shut down | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Aug 09 1996 07:32 | 8 |
|
We lost slurry and Ammonia for 45 minutes due
to a tech in 6.1 shuting down bulk chemical
distribution to the entire fab for a sink issue.
All tools came back on line fine.
D shift
|
61.47 | bolts!!!! oh nuts.... | SUBPAC::LANDRY | | Mon Aug 19 1996 20:32 | 7 |
| Bad day with bolts on the Strasbaugh carriers. Three plates have
broken screws in the back side. All happened as the carriers were
being dis-assembled. An SBA has been cut and the plates will be sent
out for repair.
A shift
|
61.48 | WW9 Semi Annual P.M. for B2 Partially Completed | ASDG::POIRIER | | Sat Aug 31 1996 20:34 | 7 |
|
CMP.B2's Semi Annual P.M.
------------------------
The Semi Annual P.M. for WW9 is almost complete. (6 items remaining)
A Partially completed check sheet has been posted in note #30.
C Shift.
|
61.49 | Slurry Delivery System Failure | ASDG::POIRIER | | Fri Sep 06 1996 20:02 | 9 |
|
Slurry Delivery Failure:
-----------------------
The Slurry pressure dropped to zero twice today. The blue warning light
on the wall did not flash.
C Shift.
|
61.50 | slurry delivery troubles | SUBPAC::LANDRY | | Sat Sep 07 1996 19:40 | 6 |
| As was the case the day prior, the slurry system went off today. We
contacted Ashland, they swapped over to a new filter and room pressure
returned to normal. We were able to stop processing wafers before any
breakage could occur.
Brad / Doc
|
61.51 | Slurry problem again | FABSIX::L_SHKOLNIKOV | | Fri Sep 13 1996 00:02 | 5 |
| The slurry system went off about 10:30 pm this night. I contacted Ashland,
so they changed filter with a new one and room pressure returned to normal.
No wafer breakage had occur during the slurry system shutdown.
Leon
|
61.52 | DI glitch | FABSIX::O_WHITFORD | | Sat Sep 14 1996 08:44 | 6 |
| At @ 06:00 this morning, we had a brief (2~3 second) hit on the DI
system. It was long enough to put CMP.B1, and CMP-Scrub.A2 & .A3 into
alarm for low/no DI pressure. Before we could get into the room, the DI
system was fine again. B1 was running break ins, B2 & B3 were in wet
idle at the time. No wafers were lost, just got the ops all charged up
on adrenalin... :)
|
61.53 | Particle Problems | ASDG::POIRIER | | Thu Sep 26 1996 20:08 | 42 |
|
Particles:
----------
CMP.A1.CMP.A2 & CMP.B2 are all DOWN for high particles. Kyle wrote
new recipies for with longer rinse cycles for the Strasbaughs. The
Particles dropped from the 4000-6000 range to the 200-300 range.
This is the same approach that was taken with the Speedfams a few
months ago when we had a similar particle problem. We are currently
waiting for the PSS to implement an EDCO with the appended recipes
before brining the Strasbaughs UP.
CMP.B2 Particles:
-----------------
The tool and carriers were thoroughly cleaned and the pad was changed.
The system failed MQC's for particles. 1500 - 2000 ave across 5 heads.
All 5 carriers are being changed and a particle check will be run.
2 carriers have had 2000+ wafer runs on them. The other 3 have had
less than 1000 wafers run.
Its not clear at this time if the particle problem with CMP.B2 and the
particle problem with the Strasbaughs are related. The particle scans
on the 6200 do resemble the patterns that were observed a few months
ago when we struggled with the big particle problem. Also, adding extra
rinse steps to the polishing recipes is having a direct cause and
effect, much like the previous particle problem. If the other Speedfams
start showing high particles at their next MQC's, I think its fair to
assume that the old particle problem is back.
Eric Bodensieck has been talking to Ashland about the timimg of filter
changes and timing of slurry barrel changes & lot #'s. He has asked
us to try mounting a point of use slurry filter on one of the effected
tools. I have ordered 10 micron and 5 micron filters. We will ask D
shift to try them tonight.
C Shift.
|
61.54 | ex | LUDWIG::RIDLON | Eliminate the oblivious | Sun Sep 29 1996 05:08 | 9 |
| FYI, when we lost slurry to the room the blue warning light on the wall
did not light up. When the cdu was checked the pump was in fact not
pumping. The cdu was alarming for drum empty but Ashland informed us
that the cdu will not stop pumping because of that. He further
explained that the reason it did stop was due to the filter being
clogged.
B-shift
|
61.55 | Removal rates on the slide... | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Oct 03 1996 07:22 | 6 |
|
We are currently seeing a decrease in removal rates on all
of the polishers since the slurry flushing. Some more obvious
than others. This issue will be addressed in the a.m. passdown.
BJ
|
61.56 | New batch of slurry on line | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Oct 04 1996 00:09 | 5 |
|
Just an FYI... Ashland just called (23:10, 10/3) to
let me know that the "new" batch of slurry is now on-line.
no... I don't mean the internet.... BJ
|
61.57 | {Power Glitch} | FABSIX::J_SWENSON | | Sat Oct 12 1996 20:05 | 8 |
| At approximately 1045 there was a power glitch throughout the Fab.
No wafers were lost. Di was lost for 5-10 minutes. The tools were
unloaded and reset, then put to wet idle when Di was restored.
No major problems resulted.
C-Shift
|
61.58 | slurry supply inadequate | SUBPAC::LANDRY | | Sun Oct 13 1996 20:21 | 8 |
| At approximately 12:15 pm today, the slurry supply to the room dropped
in pressure to a level that ALL the systems needed to halt processing.
Ashland Chemical was contacted (by pager). Dave Spang reported to the
call. He arrived to find the filter clogged. It was changed out and
the supply was in abundance. Manufacturing was back on line at
approximately 1:45 PM.
A shift
|
61.59 | slurry loop flush / change | SUBPAC::LANDRY | | Tue Oct 22 1996 20:27 | 23 |
| A rough, but productive day......
We arrived to find the entire room down due to excessive particles on
all the polisher MQC's. The particle problem increased every day
through the week. Slurry was again considered the culprit. The
previous days slurry flush was not performed properly. They did flush
the day tank, but did not flush the entire loop. So, to prove that the
slurry was the problem, we hooked up a 55 gallon drum to A1. This drum
had been delivered fresh this morning, a new lot as well. The particle
readings with this barrel were very good. So, we had Ashland dump all
the slurry from the day tank and the entire loop. They then flushed
the loop and put in the new slurry. ALL systems passed for particle
checks.
A note of interest......the slurry "lot" that was in the day tank and
loop was NOT contaminated. Rather, something within the loop had
contaminated it. This was proven by bringing a 55 gallon drum of the
same lot as the contaminated loop into the room and running directly
from the barrel to the normal input port (at the three way valve).
Particles were fine!
A shift + various helpers
|
61.60 | Slurry system flushed | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Oct 25 1996 00:49 | 8 |
|
Ashland took slurry system and flushed until ~11:15pm.
As instructed, they put a 30u Omnius filter in-line. All
of the tools came up for flow. Time will tell if the purge
helped with the particles.
BJ
|
61.61 | A slurry flush a week keeps the particles away! | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Oct 25 1996 07:15 | 7 |
|
Recap....
All tools came in less than ~150 particles.
A1, A2, B2 and B3 all set and UP/RUNNING!!!!!
|
61.62 | House Slurry System Failed - No Alarms Were Activated | ASDG::POIRIER | | Fri Oct 25 1996 20:25 | 12 |
|
House Slurry System Failed Again:
--------------------------------
Ashland called at 6:30 to say we lost slurry. As a result, two product
wafers were broken on CMP.A2. The Ashland tech informed us that he
removed the clogged 30 um filter and replaced it with a bag style filter.
Process Engineering was called at home and messages were left to
contact the CMP area ASAP. Due to this filter change, all CMP tools
will need MQC's for particles before production can resume.
C Shift.
|
61.63 | Slurry flushed.... again.... | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Oct 26 1996 05:07 | 6 |
|
Tools failed for particles. Slurry system was flushed and
a 30u osmmonic filter was installed. As of right now, 04:05,
ALL polishers have passed particles.....
D shift
|
61.64 | slurry filter changed | SUBPAC::LANDRY | | Sun Oct 27 1996 19:35 | 6 |
|
Ashalnd Chemical changed the slurry filter this morning at
about 8:30 AM. No particles problems all day!
A shift
|
61.65 | latest filter change | SUBPAC::LANDRY | | Mon Oct 28 1996 19:17 | 5 |
|
Ashalnd Chemical changed out the slurry filter at ~2:30 PM
A shift
|
61.66 | House Slurry System Purge Complete | ASDG::POIRIER | | Fri Nov 01 1996 19:38 | 13 |
|
House Slurry System:
-------------------
The system was sucesfully purged with KOH and Slurry today. B2 just
passed MQC's with 200 adders. System is as clean as its ever been.
Note: B1 will need a slurry purge of at least 1/2 hour before MQC's are
begun. We didn't close the valves in the valve box under the
floor. Since B1's slurry solenoid valves are in its multihead,
KOH has filled the lines up to the slurry valves in the
multihead. A 1/2 hour slurry purge should clean the KOH out.
C Shift.
|
61.67 | Low Removal Rates On Speedfams - Pads Suspected | ASDG::POIRIER | | Sat Nov 02 1996 19:44 | 15 |
|
Low Removal Rates On Speedfam B2 & B3:
-------------------------------------
The removal rates on B2 & B3 are now less than 1000 A. We have verified
that the Strabaughs are no longer back filling D.I. into the house
slurry system. We have taken the pad off B3 and will replace it with
one from a lot other than (6DZ-JP5). We suspect that the pads might be
the cause of the low removal rates.
FYI: We can't find a polytex pad for the 2nd table in the stockroom.
We will only be changing the primary pad at this time.
C Shift.
|
61.68 | 50k's are scarce | LUDWIG::RIDLON | Eliminate the oblivious | Sun Nov 03 1996 05:20 | 3 |
| Both A1 and B1 sat for ~8 hours waiting for 50k breakin wafers so that
MQC's could be performed. A1 had a scheduled pad change and B1 was
coming back from being down.
|
61.69 | evac | FABSIX::M_RIDLON | Eliminate the oblivious | Sun Dec 08 1996 06:17 | 6 |
| Power glitch took out solvent exhaust which in turn activated the local
evac's for fab6. The tools sat for an hour before I could come in and
put them in wet mode. No product was lost. The pads were a little dry
so they were conditioned and rinsed thuroughly and all tools placed in
wet mode for remainder of evacuation (3+hours).
|
61.70 | Polish Pad storage cabinet moved | FABSIX::R_GEE | | Tue Dec 10 1996 19:06 | 6 |
| Per engineering we moved the black pad storage cabinet
into the CMP.B3 chase.
A shift
|
61.71 | MEGASONIC PUMP | LUDWIG::WESTPHAL | | Fri Dec 20 1996 19:58 | 4 |
| The pump for the megasonic sink in the carrier rebuild room is not
working. I pulled the filter housing off and found shim in it.
suspect that there is probably one in the pump. Did not have time to
address the issue today.
|
61.72 | Carrier rebuild room megasonic sink pump | FABSIX::R_GEE | | Sun Dec 22 1996 17:11 | 6 |
| The megasonic sink in the carrier rebuild room wasn't working.
Found breaker B601 had tripped, reset the breaker. Due to the
findings on "C" shift we checked the pump and lines for any
foreign material, none found. Ran the pump for 1 hour, all set.
A shift
|
61.73 | slurry lost | SUBPAC::LANDRY | | Mon Dec 23 1996 19:31 | 8 |
| We lost slurry to the room today. Systems were halted before any
breakage happened. Oh ya, the blue strobe lights worked! Contacted
the Ashland folks, who trouble shot the problem to an over fill sensor
on the day tank. Resolved said problem and we were back on line. Took
about 10 minutes from down to up.
A shift
|
61.74 | Ammonia pressure problem resolved | SUBPAC::LANDRY | | Tue Jan 14 1997 16:26 | 22 |
|
Ammonia flow, or better yet a lack of flow, became a big issue this
morning. For several days/shifts, the tools in the room have been
alarming for low ammonia flow. Each time Ashland was contacted, they
would call back within minutes saying everything looks good on their
end. But with all the tools losing pressure at the same time, the
problem had to have one central source. Therefore, the Ashland folks
decided to camp out at the system awaiting for the problem to occur.
It didn't take to long to find the problem. Their bulk delivery unit
had a software command that would actuate a valve (which was either
purging or flushing the line) to open. Each time this valve opened,
the room supply would drop down below 5 PSI to the room. When the
valve closed, the pressure was back in seconds. This explains why they
were unable to first locate the problem. By the time the call was
placed to them, coupled with their short walk to the delivery system,
the purge loop was activated then shut off. So they would arrive to
find the proper pressure and all working fine. They have removed the
software command from the software. The problem has not re-occurred
since this removal.
A shift
|
61.75 | Lost slurry, no room alarm | FABSIX::L_SHKOLNIKOV | | Thu Jan 16 1997 07:10 | 5 |
| AT @ 06:05 this AM, we lost ALL slurry to the CMP room. We notified
supervision, who in turn paged Ashland Chemical. We have yet to hear
back from them although there are supposed to be two Ashland personnel
on site around this time of day. As it was, all tools were down for
external/facilities for @ 1 hr this morning..
|
61.76 | ACC phone #'s | STRATA::GOODWIN | | Fri Jan 17 1997 14:11 | 37 |
| From: HOTLNE::AGRESTI 17-JAN-1997 13:57:41.15
To: SCOMAN::GOODWIN
CC: AGRESTI
Subj: Ashland Contacts
Dennis:
Ashland's on-site hours are:
Mon - Fri 6:00am - midnite
Sun 3:00pm - midnite
Tel. Extension is 6112
Pagers:
On-site 508-489-4395 During hours
or 660 FN1
On-call 508-489-4393 After hours
Contacts: ACC pager DEC pager Home phone
David Spang 508-387-0797 674 fn1 508-649-3357
George Ingles 508-489-4564 665 fn1 508-368-1538
David Agresti 508-387-0924 508-541-7150
These will be all the numbers you need. We are working w/
Digital to get this down to one number (several pagers on the
same phone number) so that you don't need a list.
regards,
david
|
61.77 | DEPARTMENT QUALITY SPEC FORMS | LUDWIG::JONES | | Tue Jan 21 1997 06:25 | 210 |
| Attachment 7.1 FAB 6 RTC CHANGE FORM
===============================================================================
REQUESTOR:_______________________________________ DATE:__________________
===============================================================================
RTC ENTITY NAME:_______________________________ ADD( ) DELETE( ) CHANGE( )
(Maximum 12 characters)
DESCRIPTION:___________________________________________________________________
(Maximum 32 characters)
VENDOR:_____________________________ MODEL NUMBER:__________________________
SERIAL NUMBER:______________________ ASSET NUMBER:__________________________
INSTALL DATE:_______________________ ACCEPTANCE DATE:_______________________
LOCATION (BAY NUMBER):______________ KEY TOOL GROUP:________________________
(See DBA for list)
(Do not submit without either capital asset number or serial number)
TECHNOLOGY:
DIFF( ) FILMS( ) IMPLANT( ) PHOTO( ) METROLOGY( ) E-TEST( ) ETCH( )
===============================================================================
PM schedule, spec and checksheet is in place? YES( ) NO( )
PM spec includes Lockout/tagout requirements? YES( ) NO( )
Start-up/Shut-down procedure in PM spec? YES( ) NO( )
Spare parts are in place or on order:_________________________________________
(Appropriate stockroom planners signature)
System name tag is installed? YES( ) NO( )
System acceptance documents on file? YES( ) NO( )
"No Calibration Required" label is installed? YES( ) NO( )
NOTE: If 'NO' to any of the above, signature of CC MGR is required.
===============================================================================
"No Calibration Required" label number: NR__________
Calibration control number:____________
Calibration lab signature:_____________________________________ Date:_________
===============================================================================
APPROVALS
Equipment Engineer:_____________________________________________ Date:________
Process Engineer:_______________________________________________ Date:________
Operations Bus Mgr:_____________________________________________ Date:________
RTC DBA Signature (entry complete):_____________________________ Date:________
Attachment 7.2 DEINSTALLATION CHECKSHEET
(This form should be posted in the PM notes file by the
appropriate equipment engineer for the affected
equipment. Appended mail messages will indicate
approval by the business groups involved.)
RTC NAME OF EQUIPMENT:_________________________________________________
EQUIPMENT MANUFACTURER:________________________________________________
MODEL NUMBER:___________________ SERIAL NUMBER:______________________
ASSET TAG NUMBER:______________________________________________________
DATE SCHEDULED FOR REMOVAL FROM PRODUCTION:____________________________
=======================================================================
APPROVAL FOR REMOVING TOOL FROM COST CENTER MANAGER:
COST CENTER MANAGER:_________________________ DATE:______________
=======================================================================
TOOL PLACED ON DIAL 90 DAYS?
FINANCE APPROVAL:____________________________ DATE:______________
=======================================================================
EQUIPMENT DISPOSITION GROUP NOTIFIED 120 DAYS IN ADVANCE OF TOOLING
AVAILABILITY?
FINANCE APPROVAL:____________________________ DATE:_____________
=======================================================================
DECONTAMINATION PROCEDURE DONE?______________
EH&S APPROVAL:_____________________________________ DATE:___________
=======================================================================
SPARES PULLED FROM STOCKROOM?________________
MATERIALS GROUP APPROVAL:__________________________ DATE:__________
======================================================================
SPECIAL TOOLING DELIVERED TO EQUIPMENT DISPOSITION GROUP?_____________
FINANCE APPROVAL:____________________________________ DATE:__________
======================================================================
ATTACHMENT 7.3 EQUIPMENT MODIFICATION/UPGRADE SHEET
Date:
Engineer:
System:
Date sys requested:
Modification:
Est time duration:
Reason for mod:
Recovery for mod failure:
Mfr training required:
EEG training required:
Safety impact:
Preventive maintenance or operating
specification changes:
Documentation for this modification (provided below or identify controlled
area):
Back up systems:
Distribution:
(At minimum this distribution includes manufacturing and process
engineering supervisors and the equipment technicians within the
group.)
ATTACHMENT 7.4
Preventive Maintenance Qualification Observation Form
Equipment Name _________________________________
PM Qualification Spec Spec Qualification Trainer's
Type Number Revision Date Initials
A. _______________ __________ _______ __________ ______________
B. _______________ _______ __________ ______________
C. _______________ _______ __________ ______________
D. _______________ _______ __________ ______________
E. _______________ _______ __________ ______________
F. _______________ _______ __________ ______________
G. _______________ _______ __________ ______________
Reference Specs: Spec Number:
Working on or Near Energized Equipment 851.013
Fab 6 Protocol Specification HF60-0140
Lone Worker Policy HH20-00000-0007
Electrical Lockout/Tagout Specification HH20-00000-0011
Electrical Safety Related Work Practices HH20-00000-0020
==========================================================================
Qualification History/Comments:
__________________________________________________________________________
__________________________________________________________________________
__________________________________________________________________________
__________________________________________________________________________
Issued by:__________________________________ Badge # _______ Date _______
==========================================================================
Trainer's Signature ________________________ Badge # _______ Date _______
Trainee's Signature ________________________ Badge # _______ Date _______
Supervisor's Signature _____________________ Badge # _______ Date _______
ATTACHMENT 7.5
Preventive Maintenance Initial Qualification Approval Form
Equipment Name _________________________________
PM Qualification Spec Spec
Type Number Revision
A. _______________ __________ _______
B. _______________
C. _______________
D. _______________
E. _______________
F. _______________
G. _______________
Completion of this form enters initial qualee into the formal
documentation process and, as such, pre-empts them from going through an
offical qualification observation process.
==========================================================================
Initial Qualee Signature ___________________ Badge # ________ Date _______
|
61.78 | lost slurry | FABSIX::M_RIDLON | Eliminate the oblivious | Wed Feb 26 1997 07:45 | 5 |
| We lost slurry to the room for about 15 minutes tonight. No product was
lost. Ashland was here and took care of the problem.
B-shift
|
61.79 | slurry loss | SUBPAC::LANDRY | | Sun Mar 02 1997 19:40 | 7 |
| At approximately 11:30 this morning, we lost slurry to the room.
Contacted the Ashland folks. They found that the tank had run below
the day tanks supply tube, hence sucking up some slurry and a bunch of
air. They filled the tank and we were back in business. No wafer loss
or wafer damage was observed, thankfully.
Doc
|
61.80 | NOSLURRY ALARMS | FABSIX::B_WESTPHAL | | Fri Mar 21 1997 19:53 | 7 |
|
CMP.B - No slurry alarms on all tools. The sensors caught some slurry bubbles
when the filter was being changed. Corrected the situation and
cleared the alarms.
|
61.81 | Pad's Glazing on A and B Tools | FABSIX::J_SWENSON | | Sat Mar 29 1997 19:31 | 21 |
| PAD GLAZING
-----------
Today there was incidents of pad glazing on cmp.a2,
cmp.b2 and cmp.b3.
It seems that there has been an increase of glazing
which requires extra conditioning to prevent.
Looks like we may have some hard pads within our stock.
The glazing on B2 acually caused the tool to vibrate
during polish, the pad was then reconditioned which
brought the pad back.
Will need to keep an eye out for future glazing
C-Shift
|
61.82 | Status of tools in CMP 4-1-97 | FABSIX::R_POIRIER | | Tue Apr 01 1997 06:13 | 31 |
|
CMP.A1 Has new pad (not wet), new conditioner, and
there are two new carriers in the room ready
to be put on tool.
CMP.A2 Failed 24 hr MQC's. Took off pads and conditioner
Put on new conditioner, but left pad off so I
could put the tool in wet idle to keep the carriers
and brushes wet. The carriers only have about 270
wafers on them.
CMP.B1 Up and in wet idle.
CMP.B2 Up and in wet idle. (Weekly PM done)
CMP.B3 Up and in wet idle. (Weekly PM done)
WARNING on B3. Slurry loss detect jumpered out.
(I broke the detector when I was doing the PM,
and I couldn't find the part in SGMR)
CMP-SCRUB.A1 Up and idle
CMP-SCRUB.A3 Up and idle
CMP-SCRUB.A3 Up and idle
CMP-SCRUB.C1 Up and idle (Weekly PM done)
|
61.83 | Lot slurry on B2, B3 and CMP.A tools | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri May 16 1997 04:03 | 9 |
|
Lost slurry from Systems Chemistry unit. Ashland
tried swapping pumps because the one that was on-line
sounded funny. After the swap, we still has little pressure.
The filters were checked and they were clogged. They replaced
the filter(s) and we were good to go.
BJ
|
61.84 | Slurry Purge | FABSIX::J_SWENSON | | Fri May 16 1997 20:26 | 13 |
| Particle issues on A and B tools.
--------------------------------
- The slurry system was purged today at approximately 1430.
- A1,A2 and B3 were put on barrels and run.
- At 1710 the purge was completed and all tools put back to house
system.
- As of 1930 all tools have been running without issue.
C-Shift
|
61.85 | RECIRCULATION PUMP REPLACED | FABSIX::B_WESTPHAL | Bradley BP#661 EXT# 6610 | Sat May 17 1997 20:54 | 5 |
|
* NOTE - The recirculation motor for the metal slurry distribution unit was
changed out with a new part as requested by engineering. The old pump sets in
the metals room sink needing to be Decontaminated and we need a decon sticker.
|